What is the total volume of gas produced by the action of heat on 1.00 g of silver nitrate?
Given that the gas is collected at room temperature and pressure.
1
Expert's answer
2020-04-29T13:29:21-0400

AgNO3= Ag + NO2 + 1/2O2.

1 x 1.5 / 170 = 0.009 moles.

pV=nRT. V= nRT/p = 0.009 x 8.314 x 298 / 101.325 = 0.215750599 L.